### v2

# FairShare: Auditable Geographic Fairness for Multi-Operator LEO Spectrum Sharing

Abstract
Dynamic spectrum sharing (DSS) among multi-operator low Earth orbit (LEO) mega-constellations is essential for coexistence, yet prevailing policies focus almost exclusively on interference mitigation, leaving geographic equity largely unaddressed. This work investigates whether conventional DSS approaches inadvertently exacerbate the rural digital divide. Incorporating Keplerian orbital dynamics, inter-beam co-channel interference, and three real-world constellation geometries (Starlink, OneWeb, Kuiper), we conduct large-scale, 3GPP-compliant non-terrestrial network (NTN) simulations across 20 orbital snapshots spanning 10 minutes of satellite motion. The results uncover a stark and persistent structural bias: SNR-priority scheduling induces a $1.84\times$ mean urban–rural access disparity, with temporal fluctuations reaching $3.9\times$ during favorable interference conditions. Counter-intuitively, increasing system bandwidth amplifies rather than alleviates this gap. To remedy this, we propose FairShare, a lightweight, quota-based framework that enforces geographic fairness. FairShare not only reverses the bias, achieving an affirmative disparity ratio of $\Delta_{\text{geo}}=0.68\times$ with zero variance across all orbital snapshots and interference conditions, but also reduces scheduler runtime by 3.3%. This demonstrates that algorithmic fairness can be achieved without trading off efficiency or complexity, and that it remains invariant to physical-layer dynamics. Our work provides regulators with both a diagnostic metric for auditing fairness and a practical, enforceable mechanism for equitable spectrum governance in next-generation satellite networks.
I Introduction
Figure 1: Geographic disparity (a) and the FairShare framework (b). The Physical Problem: Rural users face a structural disadvantage through two compounding mechanisms: beam gain roll-off (as beams are centered on population-dense urban areas) and elevation-dependent path loss. The figure illustrates the latter, where rural users experience lower elevation angles ($\theta_{r}<\theta_{u}$) and longer slant ranges, contributing to the overall SNR penalty (a). The FairShare Solution: The proposed framework replaces purely competition-based allocation with geographic partitioning, enforcing specific bandwidth quotas (b). Note: the SNR gap shown is illustrative; measured values are 5–10 dB.
The rapid deployment of low Earth orbit (LEO) mega-constellations by multiple operators such as Starlink, OneWeb, and Kuiper is intensifying competition for shared Ka-band spectrum. As these systems scale to thousands of satellites, dynamic spectrum sharing (DSS) has become indispensable for their coexistence [5]. Current regulatory frameworks, including those of the International Telecommunications Union (ITU) and the Federal Communications Commission (FCC), primarily emphasize interference mitigation [13, 10]; however, they largely overlook an equally fundamental question: Do existing DSS policies provide equitable spectrum access across geographically diverse users?
Fairness is inherently geographic for non-terrestrial networks (NTN). User channel quality is directly shaped by satellite elevation geometry. Users in rural and remote areas, situated at the periphery of satellite spot beams, face a systematic geometric disadvantage: significantly lower elevation angles and longer slant ranges. This increases free-space path loss and atmospheric attenuation, resulting in a structural net signal-to-noise ratio (SNR) penalty that cannot be overcome by local environmental advantages. Unlike demographic or demand-driven disparities, this disadvantage is immutable and persists regardless of user behavior. Consequently, fairness in NTNs emerges as a unique physical-layer challenge poorly addressed by terrestrial fairness frameworks.
Classical metrics such as Jain’s index [16], the Gini coefficient, and $\alpha$-fairness [19] were developed for quasi-static terrestrial networks and fail to capture LEO dynamics, where elevation-dependent propagation dictates link quality. When channel-aware or demand-driven allocation policies are applied in this context, they systematically amplify geometric advantages, translating modest SNR differences into severe disparities in spectrum access between urban and rural users.
While prior work on LEO scheduling and NTN coexistence has focused on spectral efficiency and interference control [4, 11], geographic fairness for multi-operator spectrum sharing remains unexamined. Moreover, current 3GPP Rel-17/18/19 NTN specifications provide no fairness mechanisms for multi-operator scenarios [1], leaving a critical gap between regulatory objectives and practical allocation behavior.
This paper addresses this gap through FairShare, which serves both as a benchmarking framework and a geographic-aware allocation policy. Using 3GPP TR 38.811 compliant channel models and simulating large-scale distributed user populations, we systematically evaluate conventional DSS strategies. The results reveal that SNR-priority scheduling induces a mean urban rural access disparity of $1.84\times$, with temporal fluctuations reaching $3.9\times$ during favorable interference conditions, while demand-proportional allocation yields a $1.31\times$ gap. Critically, this disparity is shown to be policy-inherent rather than scarcity-driven and persists across all three constellation geometries evaluated.
To overcome this limitation, FairShare introduces explicit, tunable geographic quotas into the allocation process. This simple intervention enforced through a centralized coordination framework representative of emerging regulatory models (see Sec. III) not only eliminates urban bias but achieves affirmative fairness ($\Delta_{\text{geo}}=0.68\times$) with zero temporal variance, while retaining high spectral efficiency. Moreover, the quota-based design reduces computational runtime by 3.3% compared to baseline schedulers.
This work makes five key contributions: (1) an open benchmarking framework for evaluating geographic fairness in multi-operator LEO spectrum sharing under 3GPP channel models with Keplerian orbital propagation and inter-beam co-channel interference; (2) the first systematic evidence that conventional DSS policies induce extreme urban-rural disparities originating from allocation design, not spectrum scarcity, validated across 20 orbital snapshots spanning 10 minutes of satellite motion; (3) FairShare, a lightweight geographic-aware policy that guarantees equitable access while maintaining high spectral efficiency; (4) cross-constellation validation demonstrating that the structural bias persists and FairShare remains effective across Starlink, OneWeb, and Kuiper geometries; and (5) concrete, evidence-based guidance for regulators on incorporating explicit geographic fairness metrics into future NTN spectrum-sharing standards.
II Related Work
II-A DSS and Fairness in LEO Systems
Ahmed et al. [5] survey DSS mechanisms between LEO and terrestrial systems, emphasizing interference management, while Giordani et al. [11] analyze Ka/Ku-band coexistence challenges in multi-operator settings, identifying regulatory gaps regarding inter-constellation coordination. Recent AI-driven approaches such as hierarchical deep reinforcement learning [20] achieve improved throughput but optimize aggregate metrics, lacking explicit fairness guarantees or geographic considerations; the opacity of learned policies further presents challenges for regulatory auditability.
Classical fairness metrics Jain’s index [16], proportional fairness, $\alpha$-fairness [19] were designed for quasi-static terrestrial channels. Ahmad [4] demonstrates that these become unreliable under LEO dynamics, where elevation-dependent SINR fluctuations obscure systematic geographic biases. Max-min and weighted proportional fairness [8] assume homogeneous user populations without geographic structure and fail to detect location-correlated discrimination.
II-B Standardization and Regulatory Gaps
3GPP Rel-17/18/19 NTN specifications [1] define physical layer procedures, timing advance mechanisms, and mobility management for satellite access but do not address fairness in the context of multi-operator spectrum sharing. The ITU Radio Regulations provide interference coordination procedures (Article 22) but lack provisions for equitable access across geographic regions. Although fairness is not explicitly defined as a key performance indicator (KPI), the IMT-2030 framework (ITU-R M.2160) mandates ubiquitous coverage and social inclusivity, implicitly requiring geographically equitable user experience [15, 12].
Recent work on equitable access to satellite broadband and LEO frequency or orbit resources, as well as NTIA’s broadband equity initiatives, emphasize geographic disparities and the need for evidence-based equity metrics, but do not specify concrete spectrum or resource allocation mechanisms at the technical level [6]. To date, no prior work systematically evaluates geographic fairness for multi-operator NTN spectrum sharing or proposes allocation policies explicitly targeting geographic equity.
III System Model and Problem Formulation
The system under consideration captures a forward-looking regulatory scenario: a downlink multi-operator LEO NTN where $O$ satellite operators share a common spectrum pool of total bandwidth $W$ (Hz), managed by a centralized coordinator akin to the FCC’s Spectrum Access System (SAS) for the Citizen Broadband Radio Service (CBRS). [^1] While current constellations operate on separate licenses with proprietary beam management, this study models a unified dynamic spectrum access environment to evaluate the efficacy of proposed fairness regulations for future shared-spectrum frameworks. The system operates in discrete time slots $t\in\mathcal{T}$, where channel states, visibility, and beam associations evolve with satellite motion. Fig. 1 illustrates the structural relationship between geographic location and channel quality that motivates this work.
III-A Network Topology
Let $\mathcal{S}=\{1,\dots,S\}$ denote visible LEO satellites and $\mathcal{U}=\{1,\dots,U\}$ the user terminals. Each satellite $s$ forms a multibeam footprint with beam set $\mathcal{B}_{s}$, and the global beam set is $\mathcal{B}=\bigcup_{s\in\mathcal{S}}\mathcal{B}_{s}$. Each user $u$ is served by beam $b(u,t)\in\mathcal{B}$ at time $t$.
Geographic Classification: Users are classified into categories $\ell_{u}\in\{\text{urban},\text{suburban},\text{rural}\}$ based on distance from urban centers. This classification is designed to capture a hypothesized effective channel quality gradient driven by two compounding mechanisms central to our study: (1) beam gain roll-off, modeling the operational premise that satellite beams are centered on population-dense urban areas to maximize capacity, leaving rural users at beam edges with reduced antenna gain; and (2) elevation-dependent path loss and clutter variations per 3GPP TR 38.811 models [2, 18].The combined effect in our model is a systematic SNR disadvantage for rural users, allowing us to investigate fairness under a capacity-centric network planning paradigm rather than one focused solely on atmospheric propagation.
Multi-Operator Coordination: The model considers $O=3$ operators sharing spectrum through a centralized coordination mechanism, representative of emerging regulatory frameworks such as the FCC’s spectrum access system for satellite services [10]. This cooperative setting isolates the inherent geometric unfairness from competitive market behaviors. Operators submit allocation requests to a central coordinator, which applies the DSS policy to distribute spectrum across all users regardless of operator identity. Inter-operator beam conflicts are avoided through time-frequency separation.
III-B Orbital Dynamics and Constellation Model
To evaluate fairness under realistic satellite motion, we employ a Keplerian orbital propagator. Each satellite’s position in Earth-Centered Inertial (ECI) coordinates is computed from classical two-body orbital elements and transformed to Earth-Centered Earth-Fixed (ECEF) coordinates accounting for Earth’s rotation:
$\mathbf{r}_{\text{ECEF}}(t)=\mathbf{R}_{z}(\omega_{E}t)\,\mathbf{r}_{\text{ECI}}(t),$ (1)
where $\omega_{E}=7.2921\times 10^{-5}$ rad/s is Earth’s rotation rate and $\mathbf{R}_{z}(\cdot)$ is the rotation matrix about the polar axis. The simulation runs $N_{\text{snap}}=20$ orbital snapshots at $\Delta t=30$ s intervals, spanning 10 minutes of satellite motion. At each snapshot, satellite positions are re-propagated and user satellite geometry (elevation angles, slant ranges) is recomputed. Only satellites exceeding a minimum elevation angle $\theta_{\min}=10^{\circ}$ (per 3GPP TR 38.811 [2]) are considered visible.
Table I defines the three Walker-Delta constellation configurations evaluated. These span the range of current LEO filings in altitude (550–1,200 km), inclination (51.9–87.9), and constellation size (648–1,584 satellites).
TABLE I: Constellation configurations evaluated.
Constellation Planes Sats/Pl. Total Alt. Inc.
Starlink Shell 1 72 22 1,584 550 km 53.0
OneWeb Phase 1 18 36 648 1,200 km 87.9
Kuiper Shell 1 34 34 1,156 630 km 51.9
III-C Channel Model
A 3GPP TR 38.811-compliant channel model is employed, capturing elevation-dependent path loss, shadow fading, and atmospheric effects. For user $u$ at elevation angle $\theta_{u}(t)$, the channel gain is
$|h_{u}(t)|^{2}=G_{s,b}(\psi_{u})\cdot G_{u}\cdot L^{-1}(\theta_{u},d_{u})\cdot\xi_{u}(t)\cdot A_{u}(t),$ (2)
where $G_{s,b}(\psi_{u})$ is the satellite beam gain as a function of the user’s off-axis angle $\psi_{u}$ from the serving beam center (Sec. III-D), $G_{u}$ is the user terminal gain, $L(\cdot)$ is the elevation-dependent path loss, $\xi_{u}(t)$ is the log-normal shadow fading, and $A_{u}(t)$ is the atmospheric loss [18, Sec. 5.5].
Path Loss: Following 3GPP TR 38.811 (clause 6.6.2) [2]:
$L(\theta_{u},d_{u})=L_{\text{FS}}(d_{u})+L_{\text{clutter}}(\ell_{u},\theta_{u}),$ (3)
where $L_{\text{FS}}$ is the free-space path loss given by $L_{\text{FS}}=32.45+20\log_{10}(f_{c})+20\log_{10}(d_{u})$ with $f_{c}$ in MHz and $d_{u}$ in km [18]. The clutter loss $L_{\text{clutter}}$ varies with geographic type $\ell_{u}$ and elevation angle, following the statistical models in 3GPP TR 38.811 [2, Table 6.6.2-1].
Shadow Fading: Large-scale fading is modeled as $10\log_{10}(\xi_{u})\sim\mathcal{N}(0,\sigma_{\text{SF}}^{2}(\ell_{u}))$, where the standard deviation $\sigma_{\text{SF}}$ varies by geographic type following 3GPP TR 38.811 [2]  [^2] Although urban environments exhibit higher shadow fading variability ($\sigma_{\text{SF}}=8$ dB), the combined effect of beam-center positioning and favorable propagation conditions results in approximately 5–10 dB urban SNR advantage depending on deployment geometry..
The signal-to-interference-plus-noise ratio (SINR) is obtained as
$\gamma_{u}(t)=\frac{p_{s,b,u}(t)|h_{u}(t)|^{2}}{\sum_{(s^{\prime},b^{\prime})\neq(s,b)}I_{u}^{(s^{\prime},b^{\prime})}(t)+N_{0}b_{u}(t)}.$ (4)
where $p_{s,b,u}(t)$ denotes the transmit power allocated from satellite $s$ and beam $b$ to user $u$ at time $t$, $I_{u}^{(s^{\prime},b^{\prime})}(t)$ is the interference power received at user $u$ from beam $b^{\prime}$ of satellite $s^{\prime}$, $b_{u}(t)$ is the bandwidth allocated to user $u$ (Hz), and $N_{0}$ is the noise power spectral density so that $N_{0}b_{u}(t)$ is the corresponding noise power. The achievable rate is
$R_{u}(t)=b_{u}(t)\log_{2}(1+\gamma_{u}(t)).$ (5)
III-D Beam Model and Interference
Each satellite generates a hexagonal 7-beam spot-beam layout centered on its sub-satellite point. The antenna gain for user $u$ at angular offset $\psi_{u}$ from beam center follows the ITU-R S.1528 [14] parabolic approximation:
$G_{s,b}(\psi_{u})=G_{\max}-12\!\left(\frac{\psi_{u}}{\psi_{3\text{dB}}}\right)^{\!2}\quad[\text{dBi}],$ (6)
clamped at $G_{\max}-25$ dB (first sidelobe floor), where $G_{\max}=30$ dBi is the peak beam gain and $\psi_{3\text{dB}}=1.5^{\circ}$ is the half-power beamwidth for Ka-band spot beams. A 4-color frequency reuse pattern is applied across the 7 beams, assigning each beam one of four sub-bands such that adjacent beams operate on distinct frequencies.
Co-channel interference arises from beams sharing the same sub-band. The interferer set for user $u$ served by beam $b$ on frequency $f_{b}$ is
$\mathcal{I}_{u}=\{(s^{\prime},b^{\prime}):f_{b^{\prime}}=f_{b},\;(s^{\prime},b^{\prime})\neq(s,b)\}.$ (7)
The 4-color reuse significantly reduces interference by ensuring that only $\sim$25% of all beams are co-channel with the serving beam. The resulting SINR is 15–25 dB lower than the interference-free SNR, placing users in a realistic operating regime for Ka-band LEO systems.
III-E Fairness Metrics
Geographic Allocation Rate (Service Availability): The expected fraction of users in category $\ell$ receiving spectrum allocation is defined as
$\rho_{\ell}=\mathbb{E}\left[\frac{|\{u:\ell_{u}=\ell\land b_{u}>0\}|}{|\mathcal{U}_{\ell}|}\right],$ (8)
where the expectation is taken over channel realizations, and $\mathcal{U}_{\ell}=\{u:\ell_{u}=\ell\}$ denotes the set of users in geographic category $\ell$. For the assumed full-buffer traffic model, unallocated users experience a transmission outage ($R_{u}(t)=0$). Therefore, $\rho_{\ell}$ serves as a direct proxy for service availability, and $(1-\rho_{\ell})$ represents the user outage probability.
Geographic Disparity Ratio: The urban-to-rural allocation gap is expressed as
$\Delta_{\text{geo}}=\frac{\rho_{\text{urban}}}{\rho_{\text{rural}}}.$ (9)
A value of $\Delta_{\text{geo}}=1$ indicates perfect geographic fairness; $\Delta_{\text{geo}}>1$ indicates urban bias. This metric is preferred over aggregate indices because it explicitly captures the spatial structure of inequality. Unlike global metrics that can mask localized starvation, $\Delta_{\text{geo}}$ enables straightforward policy targets (e.g., mandating $\Delta_{\text{geo}}\leq 1.5\times$) directly addressing the regulatory concern of the digital divide.
Average SINR: The average signal-to-interference-plus-noise ratio of allocated users is computed as
$\bar{\gamma}=\frac{1}{|\mathcal{U}_{\text{alloc}}|}\sum_{u\in\mathcal{U}_{\text{alloc}}}\gamma_{u}.$ (10)
Jain’s Fairness Index: We compute Jain’s index for user rates,
$J=\frac{\left(\sum_{u=1}^{U}R_{u}\right)^{2}}{U\sum_{u=1}^{U}R_{u}^{2}},$ (11)
to benchmark against traditional notions of fairness. We employ this metric in Sec. VII to demonstrate that high aggregate fairness ($J\approx 1$) can paradoxically coexist with severe geographic discrimination.
III-F Problem Formulation
The spectrum allocation problem can be formulated as a constrained optimization that balances efficiency and geographic fairness:
$\max_{\{b_{u}\}}\quad$ $\sum_{u\in\mathcal{U}}b_{u}\log_{2}\!\bigl(1+\gamma_{u}\bigr)$ (12)
s.t. $\sum_{u\in\mathcal{U}}b_{u}\leq W$ (12a)
$\sum_{u\in\mathcal{U}_{\ell}}b_{u}\geq q_{\ell}\,W,\quad\forall\ell$ (12b)
$\rho_{\ell}\geq\rho_{\min},\quad\forall\ell$ (12c)
$b_{u}\geq 0,\quad\forall u$ (12d)
For notational simplicity, we drop the explicit time index and write $b_{u}$ and $\gamma_{u}$ as representative per-slot bandwidth allocation and SINR, respectively. Objective (12) maximizes sum-rate (spectral efficiency) with constraint (12a) to enforce total bandwidth, constraint (12b) to guarantee minimum geographic quotas, and constraint (12c) to ensure minimum allocation rates per region.
IV DSS Policy Design
Without constraint (12b), the optimal solution allocates exclusively to high-SNR users (urban), yielding maximum efficiency but extreme unfairness. FairShare implements (12b) through explicit geographic partitioning, then solves the remaining sum-rate maximization greedily within each region.
Conventional channel-aware policies create geographic unfairness because channel quality is structurally correlated with geographic location[7]:
  • Channel-Geography Correlation: Although urban environments introduce higher local clutter loss, urban users in this model experience systematically higher net SNR than rural users. This is because the geometric advantage of higher elevation angles (shorter slant range) dominates the additional clutter loss. This structural gap is fundamental to wide-area LEO coverage.
  • Amplification Effect: When policies select users by channel quality (Priority) or combine channel with demand (Demand Proportional), urban users are systematically favored. The multiplicative nature of Demand Proportional where urban users have both better channels and higher demand creates catastrophic unfairness.
IV-A FairShare: Geographic-Aware Allocation
FairShare is grounded in a guaranteed-minimum principle: an immutable geographic disadvantage should not systematically bar users from service. It operationalizes this by enforcing minimum geographic quotas, rejecting pure merit-based (SNR-priority) and utilitarian (throughput-maximizing) objectives that amplify structural bias. Fairness is defined at the user level, independent of operator identity, ensuring equitable access regardless of which constellation serves a given user.
FairShare guarantees geographic quotas while optimizing channel quality within each region. The key insight is that geographic fairness and spectral efficiency can coexist through a partition-then-optimize approach:
  1. 1.
    Partition: The bandwidth is divided into geographic quotas: $W_{\ell}=q_{\ell}\cdot W$ for each category $\ell\in\{\text{urban},\text{suburban},\text{rural}\}$, where $\sum_{\ell}q_{\ell}=1$.
  2. 2.
    Rank: Within each region, users are sorted by channel quality $\gamma_{u}(t)$ in descending order.
  3. 3.
    Allocate: The bandwidth is assigned to top users in each region until the quota is exhausted, guaranteeing minimum allocation rate $\rho_{\min}$.
Quota Selection: Setting $q_{\ell}$ proportional to population share of region $\ell$ with compensation for channel disadvantage achieves $\Delta_{\text{geo}}=1.0\times$. Quotas can be adjusted for affirmative access policies favoring underserved regions.
Algorithm 1 FairShare Allocation
1:Users $\mathcal{U}$, bandwidth $W$, quotas $\{q_{\ell}\}$, SINR $\{\gamma_{u}\}$, min per-user bandwidth $b_{\min}$
2:Allocation $\{b_{u}\}$
[…]
4:for each category $\ell\in\{\text{urban},\text{suburban},\text{rural}\}$ do
5:  $W_{\ell}\leftarrow q_{\ell}\cdot W$ $\triangleright$ Region bandwidth quota
6:  $\mathcal{U}_{\ell}\leftarrow\{u:\ell_{u}=\ell\}$ $\triangleright$ Users in region
7:  $n_{\ell}\leftarrow\min\!\bigl(\max(1,\lfloor N_{\text{alloc}}\cdot q_{\ell}\rfloor),\;|\mathcal{U}_{\ell}|\bigr)$ $\triangleright$ User slots for region
8:  Sort $\mathcal{U}_{\ell}$ by $\gamma_{u}$ descending
9:  for $u$ in top $n_{\ell}$ of $\mathcal{U}_{\ell}$ do
10:    $b_{u}\leftarrow W_{\ell}/n_{\ell}$
11:  end for
12:end for
13:return $\{b_{u}\}$
Complexity: FairShare runs in $O(U\log U)$ per region for SNR-based sorting plus $O(U)$ for classification. This is negligible relative to channel estimation.
IV-B Optimality Analysis
Proposition 1 (Pareto Optimality).
Under fixed geographic quotas $\{q_{\ell}\}$, FairShare achieves the maximum sum-rate among all policies satisfying the same quota constraints.
Proof.
Given quota constraints that fix the bandwidth allocated to each geographic region, any deviation from SNR-based user selection within regions strictly decreases the sum-rate without improving $\Delta_{\text{geo}}$. Since FairShare selects the highest-SNR users within each quota, it maximizes the objective (12) subject to (12b). Thus, FairShare lies on the Pareto frontier of the fairness-efficiency tradeoff. ∎
Setting $q_{\ell}$ proportional to population share yields $\Delta_{\text{geo}}\approx 1.0\times$. The default configuration ($q_{\text{rural}}=35\%$ vs. 30% population share) intentionally over-allocates to rural users, achieving affirmative fairness ($\Delta_{\text{geo}}=0.68\times$).
Design Philosophy: FairShare is intentionally simple. The partition-then-optimize approach embodies the principle of guaranteed minimum access ensuring no geographic region falls below a defined allocation threshold regardless of channel conditions. This simplicity is a feature for three reasons: (1) Regulatory Transparency: Quota-based policies are interpretable by policymakers and auditable by regulators, unlike opaque learned policies; (2) Provable Guarantees: FairShare provides deterministic fairness bounds, whereas optimization-based approaches offer only statistical guarantees; (3) Deployment Practicality: The algorithm’s low complexity ensures real-time operation on general-purpose processors without requiring specialized hardware, as evidenced by its 3.3% runtime reduction compared to priority scheduling (Sec. VI).
V Evaluation Setup
V-A Simulation Framework and Network Topology
The proposed FairShare policy is evaluated using a high-fidelity, system-level simulation framework built in TensorFlow following 3GPP TR 38.811 channel models, with graphics processing unit (GPU) acceleration. [^3] https://github.com/CLIS-WPI/FairShare We employ a multi-snapshot Monte Carlo approach incorporating Keplerian orbital propagation (Sec. III-B). The simulation runs $N_{\text{snap}}=20$ snapshots at 30-second intervals, spanning 10 minutes of satellite motion. At each snapshot, satellite positions are re-propagated and the full channel model including inter-beam co-channel interference (Sec. III-D) is recomputed. Three Walker-Delta constellation configurations are evaluated (Table I): Starlink Shell 1 (1,584 satellites at 550 km), OneWeb Phase 1 (648 satellites at 1,200 km), and Kuiper Shell 1 (1,156 satellites at 630 km) [9].
The coverage area is centered at the New York City metropolitan area ($40.7^{\circ}$N, $74.0^{\circ}$W). The network scenario models a multi-operator environment where $O=3$ operators share a common spectrum pool of bandwidth $W=300$ MHz at a carrier frequency of $f_{c}=20$ GHz (Ka-band). User association follows a Best-SINR policy, where each user terminal connects to the visible satellite offering the highest instantaneous channel quality.
V-B Baseline Policies
We implement three conventional allocation policies, which are introduced in continuation, to contextualize FairShare’s performance.
Equal Static: Uniform random allocation independent of channel quality: $b_{u}(t)=W/|\mathcal{U}_{\text{active}}|$ for randomly selected users. This provides a fairness baseline but ignores spectral efficiency.
SNR Priority: Resources are given to users with highest channel quality $\gamma_{u}(t)$. Users are ranked by instantaneous SINR, and the top fraction is allocated resources. This prioritizes users with higher elevation angles, maximizing spectral efficiency .
Demand Proportional: The resource allocation is weighted by both demand and channel quality: $s_{u}(t)=d_{u}(t)\cdot(1+\gamma_{u}(t)/\gamma_{\max})$. This policy serves as a proxy for commercial traffic-shaping strategies that prioritize high-density service areas.
V-C Geographic User Distribution and Channel Model
Figure 2: Geographic distribution of 1,000 simulated users centered at NYC ($40.7^{\circ}$N, $74.0^{\circ}$W). Urban (50%), suburban (20%), rural (30%).
Users are classified by distance from the metropolitan center (Fig. 2): Urban (50%, Gaussian $\sigma\approx 5.5$ km), Suburban (20%, 22–55 km annulus), Rural (30%, 55–165 km ring). Channel propagation follows 3GPP TR 38.811 [2] with EIRP == 45 dBW, user terminal gain == 30 dBi, noise figure == 2 dB [3, 17], and location-dependent shadow fading ($\sigma_{\text{SF}}=8$ dB urban, 4 dB rural). A full-buffer downlink traffic model ($T=100$ slots per snapshot) provides a worst-case contention baseline.
Statistical Methodology: Each policy is evaluated over $20\times 50=1{,}000$ independent samples per constellation. Standard deviations reflect temporal variability due to orbital dynamics rather than statistical uncertainty.
VI Results
VI-A Geographic Allocation
The structural SNR gap (urban users exhibit $\sim$6 dB higher median SNR than rural users due to favorable elevation geometry) directly translates to allocation bias under channel-aware policies. Table II presents the allocation performance for the Starlink Shell 1 constellation, averaged over $20\times 50=1{,}000$ samples.
Key Findings: Priority scheduling yields a $1.84\times$ mean disparity with substantial temporal fluctuations (std =0.93=0.93), peaking at $3.9\times$ during low-interference orbital configurations. It results in a mean service outage probability of $74.3\%$ for rural users ($\rho_{\text{rural}}=25.7\%$), compared to $60.7\%$ for urban users. FairShare (with 35% rural quota) reduces the rural outage probability to $59.0\%$, effectively bridging the service availability gap with zero temporal variance.
TABLE II: Allocation rates and fairness (Starlink Shell 1, $W=300$ MHz, $20\times 50$ samples, with interference). $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$; std reflects temporal variability.
Policy Urban Rate Rural Rate $\boldsymbol{\Delta_{\text{geo}}}$
Equal Static 35.2 $\pm$ 0.2% 35.2 $\pm$ 0.3% 1.01 $\pm$ 0.01
SNR Priority 39.3 $\pm$ 3.9% 25.7 $\pm$ 8.3% 1.84 $\pm$ 0.93
Demand Prop. 38.7 $\pm$ 0.2% 29.8 $\pm$ 0.3% 1.31 $\pm$ 0.02
FairShare 28.0 $\pm$ 0.0% 41.0 $\pm$ 0.0% 0.68 $\pm$ 0.00
Figure 3: Geographic disparity ($\Delta_{\text{geo}}$) over 20 orbital snapshots (Starlink Shell 1). SNR Priority exhibits large temporal fluctuations driven by interference geometry changes, while FairShare remains perfectly flat at $0.68\times$.
Temporal Variability: A striking finding, visualized in Fig. 3, is the large temporal variability of SNR Priority ($\text{std}=0.93$) compared to FairShare ($\text{std}=0.00$). As satellites orbit, the beam overlap pattern changes, causing co-channel interference to oscillate with an approximately 5-minute period. The Pearson correlation between average SINR and $\Delta_{\text{geo}}$ for Priority scheduling is $r=0.69$: when interference is low (SINR $\approx$ 28 dB), the urban–rural quality gap is amplified and disparity peaks at $3.9\times$; when interference is high (SINR $\approx$ 10 dB), it partially equalizes users, reducing disparity to $\sim$$1\times$. FairShare’s quota mechanism renders it completely invariant to these dynamics the allocation counts per region are deterministic, yielding $\Delta_{\text{geo}}=0.68\times$ at every snapshot. This demonstrates that orbital dynamics expose priority-based policies as not only unfair on average but temporally unstable, while FairShare provides guaranteed fairness.
Cross-Constellation Validation: Table III demonstrates that the structural bias is not an artifact of a specific constellation geometry. All three constellations spanning altitudes from 550 to 1,200 km and sizes from 648 to 1,584 satellites exhibit significant urban bias under Priority scheduling, with $\Delta_{\text{geo}}$ ranging from $1.60\times$ to $1.84\times$. FairShare achieves an identical $\Delta_{\text{geo}}=0.68\times$ across all three, confirming that its quota mechanism is robust to constellation design.
TABLE III: Cross-constellation comparison of geographic disparity.
$\boldsymbol{\Delta_{\text{geo}}}$
Constellation Alt. Sats Priority FairShare
Starlink Shell 1 550 km 1,584 1.84 $\pm$ 0.93 0.68 $\pm$ 0.00
OneWeb Phase 1 1,200 km 648 1.77 $\pm$ 0.59 0.68 $\pm$ 0.00
Kuiper Shell 1 630 km 1,156 1.60 $\pm$ 0.49 0.68 $\pm$ 0.00
Interference Impact: Table IV compares system behavior with and without inter-beam co-channel interference. The 4-color frequency reuse and hexagonal beam layout reduce SINR by approximately 15 25 dB relative to interference-free SNR, placing users in a realistic operating regime. Notably, interference moderately reduces the mean disparity for Priority scheduling (from $2.55\times$ to $1.84\times$) because co-channel interference partially equalizes channel conditions across regions. FairShare remains invariant ($0.68\times$) regardless of interference, confirming its robustness.
TABLE IV: Impact of inter-beam co-channel interference (Starlink Shell 1). Average SINR is computed over allocated users only.
Avg. SINR (dB) $\boldsymbol{\Delta_{\text{geo}}}$
Policy No Intf. With Intf. No Intf. With Intf.
Equal Static 42.3 20.8 1.00 1.01
SNR Priority 47.2 32.7 2.55 1.84
Demand Prop. 42.9 22.0 1.36 1.31
FairShare 46.6 32.1 0.68 0.68
Bandwidth Sensitivity: Table V(a) reveals that disparity under Priority scheduling increases with bandwidth. At 200 MHz, the structural bias emerges ($1.29\times$); at 300 MHz it reaches $1.65\times$. Below 200 MHz, the system is saturation-limited and all policies yield $\Delta_{\text{geo}}=1.00$. FairShare maintains $0.72\times$ regardless of bandwidth.
Quota Tunability: Table V(b) demonstrates FairShare’s flexibility: adjusting the rural quota from 25% to 40% shifts $\Delta_{\text{geo}}$ from urban-biased ($1.11\times$) to strongly rural-favoring ($0.50\times$), enabling regulators to set precise fairness targets.
TABLE V: Sensitivity analysis (single-snapshot, no-interference baseline). $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$ ($\times$); consistent with multi-snapshot results (Table II).
(a) Bandwidth Impact
BW Equal Priority Demand FairShare
200 MHz 1.00 1.29 1.25 0.72
300 MHz 1.00 1.65 1.40 0.72
(b) Rural Quota Tunability
Quota $\boldsymbol{\Delta_{\text{geo}}}$ Interpretation
25% 1.11 Urban Bias
30% 0.80 Mild Rural Favor
35% (Def.) 0.72 Affirmative Action
40% 0.50 Strong Rural Priority
Achieves $\Delta_{\text{geo}}=0.68$ under multi-snapshot with interference (Table II).
VI-B Computational Efficiency
Benchmarked on an H100 GPU, FairShare achieves a 3.3% speedup over Priority scheduling (9.89 s vs. 10.22 s per cycle). This gain arises from the divide-and-conquer nature of quota-based allocation: sorting three smaller regional subsets is inherently cheaper than sorting one global user pool.
VII Discussion
VII-A Aggregate Metrics Mask Geographic Bias
Although all policies achieve Jain’s index $J>0.95$, SNR-priority scheduling yields $1.84\times$ mean urban-rural disparity up to $3.9\times$ during specific orbital configurations because aggregate metrics are inflated by the dominant urban population (50%). The interference-modulated oscillation in $\Delta_{\text{geo}}$ (Fig. 3) further compounds this: a regulator auditing at a single instant may observe values from $0.9\times$ to $3.9\times$. FairShare’s quota mechanism eliminates both problems, providing $\Delta_{\text{geo}}=0.68\times$ with zero variance at every snapshot. These findings underscore that geographic-specific metrics ($\rho_{\ell}$, $\Delta_{\text{geo}}$) are essential for meaningful NTN fairness auditing.
VII-B Unfairness Is Policy Inherent
The cross-constellation results (Table III) demonstrate that unfairness persists despite fundamentally different orbital parameters, and the bandwidth analysis (Table V(a)) shows disparity grows with capacity proving the bias is structural, not scarcity-driven. As Proposition 1 establishes, FairShare achieves the Pareto-optimal sum-rate within any quota constraint. Deploying more satellites or spectrum alone will not close the digital divide; regulatory interventions must prioritize fair allocation policies.
VII-C FairShare as a Regulatory Tool
The partition-then-optimize design offers regulators transparency (auditable quotas), efficiency (no throughput penalty, 3.3% runtime reduction), and practicality (real-time operation). FairShare is architecturally compatible with existing spectrum governance: in the FCC’s SAS/CBRS framework, a centralized coordinator already enforces tiered access policies. FairShare requires only a geographic classification layer and configurable quota parameters $\{q_{\ell}\}$ no modifications to operators’ PHY or MAC layers. The $\Delta_{\text{geo}}$ metric provides a standardized, time-invariant audit trail aligned with NTIA’s broadband equity mandates [6]. As the FCC extends shared-spectrum frameworks to NGSO services [10], FairShare offers a ready-to-deploy fairness module complementing existing interference management. We propose mandating the reporting of $\rho_{\ell}$ and $\Delta_{\text{geo}}$ and adopting tunable quota frameworks to set enforceable targets (e.g., $\Delta_{\text{geo}}\leq 1.2\times$).
VII-D Scope and Future Work
The Keplerian propagator suffices for our 10-minute window; higher-order perturbations matter only for longer studies. The full-buffer traffic and NYC-centric models provide conservative baselines; extension to bursty traffic, continental coverage, and proportional-fair intra-region scheduling are natural next steps. FairShare assumes a trusted coordinator; designing incentive-compatible mechanisms for competitive markets and validating with real LEO measurements remain open.
VIII Conclusion
Through 3GPP-compliant simulations with Keplerian orbital dynamics, inter-beam interference, and three constellation geometries, we have shown that SNR-priority scheduling creates a $1.84\times$ mean urban-rural disparity (peak $3.9\times$) a bias that is policy-inherent, persists across constellations, and worsens with increasing bandwidth. FairShare, a lightweight quota-based framework, achieves $\Delta_{\text{geo}}=0.68\times$ with zero variance across all conditions while reducing runtime by 3.3%. Its complete invariance to physical-layer dynamics makes it uniquely suited for regulatory enforcement, transforming a scheduling problem into actionable spectrum policy for closing the digital divide.
Acknowledgment
The work by Natanzi and Tang was supported in part by NTIA Award No. 51-60-IF007. The work by Mohammadi and Marojevic was supported in part by NSF Award 2332661.